### v3

# WikiDBGraph: A Data Management Benchmark Suite for Collaborative Learning over Database Silos thanks: *Equal Contribution

Abstract
Relational databases are often fragmented across organizations, creating data silos that hinder distributed data management and mining. Collaborative learning (CL)—techniques that enable multiple parties to train models jointly without sharing raw data—offers a principled approach to this challenge. However, existing CL frameworks (e.g., federated and split learning) remain limited in real-world deployments. Current CL benchmarks and algorithms primarily target the learning step under assumptions of isolated, aligned, and joinable databases, and they typically neglect the end-to-end data management pipeline, especially preprocessing steps such as table joins and data alignment. In contrast, our analysis of the real-world corpus WikiDBs shows that databases are interconnected, unaligned, and sometimes unjoinable, exposing a significant gap between CL algorithm design and practical deployment. To close this evaluation gap, we build WikiDBGraph, a large-scale dataset constructed from 100,000 real-world relational databases linked by 17 million weighted edges. Each node (database) and edge (relationship) is annotated with 13 and 12 properties, respectively, capturing a hybrid of instance- and feature-level overlap across databases. Experiments on WikiDBGraph demonstrate both the effectiveness and limitations of existing CL methods under realistic conditions, highlighting previously overlooked gaps in managing real-world data silos and pointing to concrete directions for practical deployment of collaborative learning systems.
I Introduction
Relational databases are prevalent in real-world applications [1, 2]. These databases are often heterogeneously distributed across clients, creating data silos [3, 4] that hinder large-scale data management. To overcome this limitation, collaborative learning (CL)—a data management paradigm in which multiple parties jointly train models on decentralized data without sharing raw data—has emerged as a promising solution [5, 6, 7, 8].
Existing CL frameworks predominantly focus on the training stage, including federated learning [4], transfer learning [9], and split learning [10]. Although CL attains strong performance on established benchmarks, including LEAF [11], OARF [12], VertiBench [13], and FedML [14], production-ready CL systems remain scarce [15], revealing a significant gap between current CL benchmarks and practical, end-to-end data management systems.
The limitations of existing CL benchmarks are illustrated by a real example from WikiDBs [16], a recently released large-scale database corpus: a collaboration between a National Monuments database and a Historic Places Registry (Fig. 1). Standard benchmarks typically approximate such a setting by perfectly partitioning a single table, implicitly assuming ideal schema and instance alignment. In practice, independently curated web databases exhibit structured yet imperfect overlap: some databases share identifiable attributes such as Latitude and Longitude, but differ in schema (e.g., ArchitecturalStyle vs. EstablishmentDate) and overlap only partially at the instance level—some monuments are historic places, but not all. As detailed later in our hybrid-overlap case study (Section VII-D), this specific component bridges two horizontally aligned clusters, enabling practical collaborative tasks such as predicting national-treasure categories for historic places using these partially vertical connections. This example motivates web-based data for CL evaluation: decentralized curation (e.g., via Wikidata) naturally produces autonomous, weakly aligned databases that more faithfully proxy real-world enterprise silos, thereby exposing essential data-management challenges—such as schema matching and fuzzy joining—that synthetic, perfectly aligned benchmarks [4] typically omit.
Figure 1: An example of collaborative learning
By investigating WikiDBs [16], we identify three common yet impractical assumptions in existing CL benchmarks and algorithm design: (1) isolation: each database is treated as an isolated data source, and connections between clients are ignored; (2) alignment: databases are assumed to be aligned either horizontally (disjoint samples, same features) [17] or vertically (disjoint features, same samples) [18]; and (3) joinable: databases can be fully joined into a single database.
Real-world data in WikiDBs [16] shows the opposite of these assumptions: (1) interconnected: some databases are interconnected according to their schema embeddings while others are not, forming a graph structure; (2) unaligned: databases are rarely aligned horizontally or vertically and are only functionally dependent on each other; even when some databases share the same set of columns, identifying column correspondences is non-trivial due to noisy names and orders; and (3) unjoinable: some databases cannot be fully joined into a single table because of the significant size, which disables most collaborative learning algorithms.
To bridge these gaps and better represent real-world CL tasks, we revisit WikiDBs [16]. Although it already extracts many relational databases from Wikidata, two challenges impede its direct use for CL: (1) explicit correlations between databases are scarce—only thousands of unconnected database pairs among 100,000 databases—which restricts the amount of data available for CL; and (2) database alignment and joinability are not provided, despite being necessary for CL.
We address these challenges by constructing WikiDBGraph, a large-scale, open-source graph of relational databases built from Wikidata [19]. To mitigate the limited explicit correlations, we leverage known relations within WikiDBs [16] and employ contrastive learning to train a model that predicts similarity between databases. The resulting similarity scores induce a graph whose nodes represent databases and whose weighted edges quantify their correlation. To address unaligned and unjoinable databases, we enrich the graph by defining and computing a set of properties for both nodes and edges, derived from database structure, schemas, and data distributions. Using these properties, we propose an automated data-mining pipeline for CL. While this pipeline yields average performance improvements over training on isolated databases, it reveals a critical performance gap compared to ideal centralized training. The contributions of this work include:
  1. 1.
    We construct WikiDBGraph, a large-scale graph of 100,000 interconnected relational databases, annotated with 13 node properties and 12 edge properties derived from database structure, to serve as a real-world benchmark for collaborative learning.
  2. 2.
    We design an automated CL pipeline and evaluate the overall improvement of mainstream CL algorithms on WikiDBGraph-correlated databases, validating the utility of WikiDBGraph.
  3. 3.
    We conduct case studies on feature-, instance-, and hybrid-overlapped database pairs, summarizing key challenges to guide future research.
WikiDBGraph dataset [20] and all codes [21] for its construction and evaluation are publicly available.
The paper is organized as follows. Section II introduces background and motivation of WikiDBGraph; Section III surveys database corpora and CL datasets; Section IV details the graph construction method; Section V analyzes graph structure and node/edge properties; Section VI evaluates the benchmark utility and downstream task performance; Section VII presents feature-, instance-, and hybrid-overlap CL case studies. Section VIII concludes the paper. For more details, this technical report also includes the extension to GitTables [22] (Section IX), the performance relation to database size (Section X), findings and open challenges (Section XI), and future directions (Section XII).
II Background
Collaborative learning (CL) trains models across clients without sharing raw data, spanning federated [5, 6, 8, 7, 23, 24], transfer [9], and split learning [10, 25, 26]. Federated learning iteratively aggregates client updates [27], using methods like FedProx [28] and FedOV [29] to mitigate data heterogeneity. When a single global model is insufficient, personalized federated learning [30, 31] tailors models to individual clients. For instance, applying SFL [30] to Fig. 1 allows the National Monuments database to personalize its model by aggregating messages from connected registries over the correlation graph.
Beyond federated learning, split learning [10] achieves collaboration by partitioning a model across clients (and/or a server) and training it through the exchange of intermediate representations and gradients; inference likewise requires the participating parties. This paradigm is particularly well-suited to feature-partitioned settings, where different clients hold complementary feature subsets. Transfer learning [9] provides another collaborative mechanism, often enabling one-shot or low-interaction knowledge transfer: one client trains a model and shares it with another client for adaptation via fine-tuning. While effective in certain communication-constrained scenarios, such one-way transfer typically offers weaker mutual adaptation than multi-round protocols such as federated learning or split learning.
Existing CL systems such as FATE [32] have unified multiple privacy-preserving paradigms—including FL, split learning, and transfer learning—and support diverse models, including linear regression, gradient-boosting decision trees, and support vector machines. Despite this progress, real-world CL deployments remain limited, largely because most algorithms target the model-training stage rather than the end-to-end data management pipeline.
III Related Work
This section reviews literature relevant to WikiDBGraph. First, we contrast it with established collaborative learning datasets and benchmarks to underscore its unique capability for modeling complex inter-client relationships, as summarized in Table I. Second, we compare WikiDBGraph with existing table and database corpora to highlight its structural novelty, with a summary provided in Table II.
Collaborative Learning Benchmarks
Existing collaborative learning (CL) benchmarks for federated learning (FL) and split learning (SL) often inadequately capture real-world, database-centric collaboration. Many are synthetic, formed by partitioning a centralized dataset to simulate distribution (e.g., FedNoisy [33], FPLLib [34], MarsFL [35], NIIDBench [4], SLPerf [36], VertiBench [13]), and thus may not reflect natural heterogeneity. Others are built from real-world sources (e.g., LEAF [11], ORAF [12], FedML [14], NUS-WIDE [37]) but assume perfect alignment into horizontal CL (disjoint samples, same features) or vertical CL (disjoint features, same samples) [17], whereas practical collaboration often involves partial overlap in both samples and features. Moreover, most benchmarks (i) model each client as a single table and (ii) implicitly treat all inter-client relationships as equally relevant, which is unrealistic for multi-table databases and heterogeneous cross-client correlations. Built from real-world Wikidata, WikiDBGraph is designed to capture these practical complexities by supporting hybrid relationships and varying degrees of correlation across clients.
TABLE I: Comparison with existing collaborative learning datasets and benchmarks
Dataset Data Partition1 Alignment across Clients2 Clients’ Data Format3 Clients’ Relationship4
FedNoisy [33] Synthetic Horizontal Table
FPLLib [34] Horizontal Table
MarsFL [35] Horizontal Table
NIIDBench [4] Horizontal Table
SLPerf [36] Horizontal Table
VertiBench [13] Vertical Table
VFLAIR [38] Vertical Table
LEAF [11] Real Horizontal Table
NUS-WIDE [37] Vertical Table
ORAF [12] Horizontal or Vertical Table
FedML [14] Horizontal or Vertical Table
WikiDBGraph Real Horizontal and Vertical Database
1How is the data distributed across clients? Synthetic: manually split from centralized dataset; Real: naturally distributed across clients.
2How is the data aligned across clients? Horizontal: disjoint samples, same features; Vertical: disjoint features, same samples.
3What is the format of the clients’ data? Table: Client data is represented as individual tables; Database: Client data is represented as relational databases.
4Whether there are explicit relationships between clients.
TABLE II: Comparison of existing real-world table or database (DB) corpora with WikiDBGraph.
Dataset Category DB Corr.1 #DBs2 #Tables2
SQLShare [39] Data-Only - 3.9k
GitTables [22] - 1M
SchemaDB [40] Schema-Only 2.5k -
GitSchemas [41] 50k 300k
BIRD [42] Schema-Data 95 694
CTU Prague [43] 83 813
Spider [44] 200 1020
SchemaPile [45] 75.6k 347k
WikiDBs [16] 100k 1.6M
WikiDBGraph (ours) Schema-Data 100k 1.6M
1The correlation between databases.
2Only the number of databases or tables with available data are counted.
Database and Table Corpora
Early tabular corpora primarily consist of isolated tables (e.g., GitTables [22]), which do not capture the relational structure inherent to real-world databases. To better reflect practice, database corpora with multiple interrelated tables have emerged, spanning schema-only resources (e.g., SchemaDB [40], GitSchemas [41]), data-only resources (e.g., SQLShare [39]), and schema-data corpora. However, existing schema-data corpora are often small-scale (e.g., the CTU Prague Relational Learning Repository [43]) or tailored to specific tasks (e.g., Text-to-SQL benchmarks such as Spider [44] and BIRD [42]), limiting their utility for general-purpose representation learning.Despite their scale, recent schema-data corpora such as SchemaPile [45] and WikiDBs [16] largely lack explicit relationships between databases, which hinders studying cross-database collaborative learning. WikiDBGraph addresses this gap by systematically constructing inter-database connections, enabling large-scale CL across distinct databases.
IV Dataset Construction
This section details the construction of WikiDBGraph (Fig. 2). We define the database relationship identification task (Section IV-A), present our approach (Section IV-B), evaluate the embedding model (Section IV-C), and describe the final graph construction (Section IV-D).
IV-A Problem Definition
This work leverages WikiDBs [16], a large-scale collection of relational databases extracted from Wikidata. We aim to construct WikiDBGraph by identifying and establishing correlations between these databases. The WikiDBs construction methodology is explicitly designed to create databases that are thematically coherent, with each one generated from a single starting topic from Wikidata. This makes the topic identifier the most direct and high-confidence signal for semantic linkage. We therefore initially define two databases as correlated if they share the same value for this foundational attribute, wikidata_topic_item_id, referred to as TID. This approach hypothesizes that two databases seeded from the exact same Wikidata topic are correlated. However, this explicit linkage based on TID identifies only 8,816 correlated pairs among 9,895 distinct databases. This results in a very limited set of connections, leaving over 90% of the databases without this explicit TID-based link. Such sparsity likely arises because many topics, though intrinsically related, have distinct TIDs within Wikidata.
The primary objective of this work is to train a model capable of uncovering these implicit correlations. Let $\mathcal{D}=\{D_{1},D_{2},\dots,D_{N}\}$ and $\mathcal{S}=\{s_{1},s_{2},\dots,s_{N}\}$ represent the set of all $N$ databases and $N$ schemas within WikiDBs, respectively, and let
$\mathcal{P}_{\text{explicit}}=\left\{(D_{i},D_{j})|\mathrm{TID}(D_{i})=\mathrm{TID}(D_{j}),i\neq j\right\}$ (1)
denote the limited set of explicitly correlated database pairs identified through shared TIDs. We aim to train a function $f:\mathcal{D}\times\mathcal{D}\rightarrow[0,1]$, where $f(D_{i},D_{j})$ outputs the predicted probability of a correlation between databases $D_{i}$ and $D_{j}$. We model this as a semi-supervised learning problem with limited positive labels in $\mathcal{P}_{\text{explicit}}$.
IV-B Approach
Figure 2: The overview of WikiDBGraph construction process
Database Serialization
To mitigate the verbosity of the original JSON files and reduce input size for subsequent processing, we serialize each $s_{i}$ and samples in $D_{i}$ into a concise textual format, denoted as abstract $t_{i}\in\mathcal{T}$. This serialization retains both structural information and an abstract of the data content. Specifically, for each database, we preserve its name, the names of its tables, and for each table, the names of its columns along with $\phi$ representative sample values (default $\phi=3$) from each column. An illustrative example of our serialization format is presented below:
Database: <database_name>
Table: <table_name_1>
- Column: <column_name_1> ; Samples: <value_1> | <value_2>
- Column: <column_name_2> ; Samples: <value_1> | <value_2>
Table: <table_name_2>
- Column: <column_name_1> ; Samples: <value_1> | <value_2>
- Column: <column_name_2> ; Samples: <value_1> | <value_2>
While sample values are included to offer a qualitative indication of the data, we do not incorporate full data distributions or the entirety of the raw data. This approach is adopted for two primary reasons: 1) schema-level information (database, table, and column names) combined with data samples is often more directly indicative of the database’s topic than comprehensive statistical distributions of all values. 2) Processing and representing the complete data for all databases would be computationally expensive and lead to excessively large serialized representations. While our manual inspection confirms that cryptic metadata (e.g., column_1) is rare in WikiDBGraph due to upstream curation [16], we acknowledge this remains a challenge in broader settings. In such scenarios, our graph formulation offers a natural mitigation: the semantics of opaque schemas can be inferred from the context of well-defined neighboring databases. We leave the empirical evaluation of this resilience for future work.
Training
We employ a contrastive learning [46] framework to train an embedding model, $f_{\theta}(\cdot):\mathcal{T}\rightarrow\mathbb{R}^{d}$, which maps serialized database schema from the input space $\mathcal{T}$ into a $d$-dimensional vector space. The distinction between positive and negative pairs, crucial for the contrastive loss, is determined by TIDs: a pair of database abstracts $(t_{i},t_{j})$ is a positive pair if they share the same TID; otherwise, it is a negative pair. The sets of positive and negative pairs are partitioned into training, validation, and test subsets in a 7:1:2 ratio, ensuring no database overlap between partitions.
Specifically, the embedding model $f_{\theta}(\cdot)$ is initialized using parameters from the pretrained encoder-only language model BGE-M3 [47]. We select BGE-M3 for its high efficiency, strong performance, and widespread adoption in semantic retrieval tasks [48, 49, 50], making it well-suited for identifying topical relationships between databases from their textual representations. The model is subsequently fine-tuned during our training process. To construct training instances, each positive pair $(t_{a},t_{b})$ from the training set is utilized, where $t_{a}$ serves as the anchor and $t_{b}$ as the positive abstract. For each anchor $t_{a}$, we sample $k$ negative abstracts that have distinct TIDs from $t_{a}$, denoted as $\{t_{n_{j}}\}_{j=1}^{k}$, from the training set. Each training instance consists of a triplet $(t_{a},t_{b},\{t_{n_{j}}\}_{j=1}^{k})$. We set the number of negative samples $k$ to 6, a value yielding sufficiently high AUC in Figure 3(a).
The model parameters $\theta$ are optimized by minimizing the InfoNCE loss. Initially, the anchor, positive, and negative abstracts are transformed into their respective embeddings:
$e_{a}=f_{\theta}(t_{a}),\quad e_{b}=f_{\theta}(t_{b}),\quad e_{n_{j}}=f_{\theta}(t_{n_{j}})\text{ for }j=1,\ldots,k.$ (2)
These embeddings, $e_{a},e_{b},e_{n_{j}}\in\mathbb{R}^{d}$, are then employed to compute the InfoNCE loss function [46] as follows:
$\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp\left(\frac{\mathrm{sim}(e_{a},e_{b})}{T}\right)}{\exp\left(\frac{\mathrm{sim}(e_{a},e_{b})}{T}\right)+\sum_{j=1}^{k}\exp\left(\frac{\mathrm{sim}(e_{a},e_{n_{j}})}{T}\right)},$ (3)
where $\mathrm{sim}(\cdot,\cdot)$ represents cosine similarity between two embedding vectors, and $T$ is a temperature hyperparameter controlling the sharpness of the distribution. The optimal embedding model $\theta^{*}$ is obtained by optimizing the InfoNCE loss function.
$\theta^{*}=\arg\min_{\theta}\mathcal{L}_{\text{InfoNCE}}\left(\theta;\{(t_{a},t_{b},\{t_{n_{j}}\}_{j=1}^{k})\}_{a,b,j=1}^{N}\right).$ (4)
(a) Test set ROC curve
(b) Distribution of similarities
(c) Database HDBSCAN clustering
Figure 3: Performance evaluation of the embedding model
(a) Node Degree
(b) Connected Component Size
(c) Community Size
Figure 4: Distribution of Graph Properties in WikiDBGraph
IV-C Evaluation of Embedding Model
We analyze the similarity distributions in Fig. 3(b) to validate the reliability of our learned correlations. The distribution of explicit TID-matched pairs is heavily skewed toward $1.0$, confirming that TIDs serve as high-precision labels with negligible false positives. Conversely, the false negatives are very prevalent according to the global distribution revealing a massive tail of high-similarity pairs lacking TID links. Consequently, leveraging these high-precision (low–false-positive) labels, our model can safely expand beyond TID links to uncover a large set of implicit correlations that TID matching misses, substantially enlarging the database network. Importantly, while these newly discovered correlations are numerous, they remain a small fraction of all possible pairs. Consequently, their discovery does not substantially bias the negative sampling strategy used during training.
The clustering of the database embeddings, projected to two dimensions using t-SNE and subsequently processed with HDBSCAN, is illustrated in Fig. 3(c). The visualization yields 11 distinct clusters (excluding “unknown”) of varying sizes that correspond to topical categories. The pronounced size imbalance—with Cluster 9 (biomedical) substantially larger than others, e.g., sports (Cluster 10)—is not a clustering artifact but reflects the long-tailed topical distribution of the source corpus, WikiDBs [16]. WikiDBs contains many biomedical databases (e.g., genes, proteins, diseases), likely due to the prevalence of tabular data in the biomedical domain.
This learned embedding model is not universal across all database corpora. However, the underlying training paradigm readily extends to other corpora, such as GitTables [22] without pre-defined labels (TIDs). We demonstrate this extension in the technical report [51].
IV-D Graph Construction
Upon obtaining the pretrained embedding model, denoted as $f_{\theta^{*}}$, the embedding vector for each database in the corpus can be derived. To construct the database graph, we perform nearest-neighbor retrieval in the embedding space and compute cosine similarities between each database embedding and its nearest neighbors. We use the faiss library [52] to enable efficient large-scale nearest-neighbor search.
Threshold methodology. We treat $\tau$ as a graph sparsification parameter rather than a tuned hyperparameter. We first obtain a data-driven baseline $\tau_{0}$ from the embedding model’s similarity predictions, using the threshold that maximizes Youden’s index. As shown in Fig. 6, this choice can produce a very dense graph, which may be undesirable for applications that require efficient construction and analysis. Accordingly, we release (i) all graph-agnostic node/edge properties computed over edges with similarity $\geq\tau_{0}$, and (ii) the complete graph-construction pipeline, so that users can instantiate graphs under any threshold suited to their budget and use case. In the paper, we additionally report three representative sparsity regimes (low/medium/high) to illustrate how key graph statistics vary with edge pruning (Table III).
Downstream evaluation and pair selection. Our downstream experiments (Section VI) focus on the highest-similarity pairs because this aligns with how most CL methods are intended to be used: performance is meaningful and comparable only when the paired parties exhibit strong semantic relatedness (e.g., substantial feature/instance overlap). Accordingly, we evaluate on a fixed set of top-ranked pairs ($\tau>0.98$) whose similarities are uniformly very high, making the downstream results effectively insensitive to the particular global sparsification threshold used for graph construction; $\tau$ mainly affects density-related graph analyses and the exploration of more challenging weak-overlap regimes.
V Dataset Details
This section introduces the structural details of the WikiDBGraph dataset (Section V-A) and subsequently describes the properties defined for the nodes and edges within this graph (Section V-B), from which three characteristics emerge.
V-A Graph Structure
The structural characteristics of the generated database graphs, as detailed in Table III and further visualized by the distributions presented in Figs. 4(a), 4(b), and 4(c), reveal the first characteristic of data in CL tasks.
Characteristic 1: Client databases form an interconnected graph with non-uniform sparsity and correlation.
The mean degree reported in Table III is in the hundreds, indicating extensive interconnection among databases. This is further supported by Fig. 4(a), which shows a substantial share of high-degree databases. However, the network is heterogeneous: the degree distribution is long-tailed—only a small fraction of databases are high-degree, whereas many reside in small clusters or communities, as evidenced by the connected-component size distribution in Fig. 4(b) and the community-size distribution in Fig. 4(c). This imbalance is further visualized in Fig. 3(c), which reveals many small clusters. Such long-tailed structure with a few high-degree “hubs” is common in real-world networks such as social networks [53] and webpage networks [54]. Collectively, these results show that the databases in WikiDBGraph are broadly interconnected, yet connectivity varies across databases, underscoring the necessity of a graph structure for CL studies.
V-B Node and Edge Properties
To further characterize data in CL tasks, we compute properties for both nodes (databases) and edges (relationships between databases) in the graph, grouped into structural, semantic, and statistical categories. Structural properties capture architectural characteristics (e.g., table counts and schema similarity). Semantic properties encode conceptual meaning (e.g., topic embeddings and their cosine similarity). Statistical properties quantify the underlying data (e.g., column cardinality and distribution divergence). Tables IV and V summarize these properties; specific details follow.
Nodes
For nodes, which represent individual databases, these properties provide a multifaceted profile that spans structure, semantics, and statistics. Structural properties define the database architecture: basic scale metrics include the total number of tables (#Tables) and columns (#Columns); CategoricalRatio indicates the proportion of columns containing non-numeric data; and ForeignKeyDensity and AverageConnection quantify internal schema complexity by measuring, respectively, the ratio of foreign keys to columns and the average number of foreign-key links per table. Semantic properties capture conceptual identity: each node is characterized by a high-dimensional DatabaseEmbed vector representing its content, its assignment to a broader topic ClusterID, and its CommunityID derived from its position within the graph structure. Statistical properties summarize the stored data: DataVolume is the total file size; AllJoinSize estimates the row count resulting from joining all tables; AverageCardinality reflects the average number of unique values per column; AverageSparsity indicates the prevalence of NULL or empty values; and AverageEntropy measures the average information content of the columns.
Edges
Complementing these node-level descriptors, edge properties quantify the nature and strength of relationships between database pairs across structural, semantic, and statistical dimensions. Structural properties measure schema similarity via JaccardTable, JaccardColumn, and JaccardType (overlaps of table names, column names, and data types), along with counts of CommonTables, CommonColumns, and CommonDataTypes; HellingerDistanceType captures dissimilarity between data-type distributions, and graph edit distance (GED) quantifies structural differences between schema graphs. Semantic properties focus on conceptual likeness: EmbedSim is the cosine similarity between database embeddings, and SimilarityConfidence reflects the certainty of this similarity prediction. Statistical properties assess correspondence in shared columns: KLDivergence measures divergence between their data distributions, while OverlapRatio quantifies the proportion of identical values.
These statistics reveal two additional characteristics of data in CL tasks.
Characteristic 2: The alignment between databases is a hybrid of horizontal and vertical alignment.
Table V indicates limited feature overlap across databases: the ratios of table and column overlaps are low on average (JaccardTable mean $0.12$; JaccardColumn mean $0.36$), and absolute overlaps remain modest (CommonTables mean $1.85$; CommonColumns mean $74$). Meanwhile, the instance overlap is also limited: the OverlapRatio averages $0.22$, indicating that shared columns contain relatively few identical values across databases. Taken together, these observations imply that most cases cannot be idealized as purely horizontal (same features, disjoint samples) or purely vertical (same samples, disjoint features); rather, the alignment across databases is partial and hybrid.
Characteristic 3: Many databases are unjoinable due to the size of the data.
Although the mean DataVolume per database is small (1.5 MB; Table IV), the average AllJoinSize, which is the estimated row count from joining all tables within a database, is $9\times 10^{16}$. Such joins are computationally prohibitive in both time and memory, rendering CL methods that assume a single, pre-joined table impractical in this setting. Consequently, practical CL must operate over relational structures directly, avoiding explicit full joins while leveraging graph-aware strategies.
TABLE III: Summary of Graph Statistical Properties ($C$: Connected Component)
Confidence1 $\tau$ #Nodes #Edges #$C$ #Isolated Nodes Degree Size of Connected Component
Min Max Mean Median Min Max Mean Median
10% 0.9252 100,000 26,879,058 64,417 58,010 0 5,872 268.79 0 1 12,054 1.55 1
15% 0.9436 100,000 17,964,868 71,235 65,126 0 4,803 179.65 0 1 10,703 1.40 1
20% 0.9555 100,000 6,197,746 78,359 73,300 0 4,527 123.95 0 1 7,960 1.28 1
1The ratio of $\tau$ increased by $\tau_{0}$; larger ratios maintains higher confidence of relationship but retain fewer edges.
TABLE IV: Summary of Node (Database) Properties in WikiDBGraph
Category Property Min Max Mean Median Description
Structural #Tables 2 100 16.11 3.0 Number of tables in the database
#Columns 7 7,955 849.32 78.0 sum of #columns across all tables
CategoricalRatio 0.4 1.0 0.92 0.92 Proportion of categorical columns
ForeignKeyDensity 0.0 0.36 0.05 0.04 #foreign keys / #columns
AverageConnection 1.0 13.5 2.71 1.50 Average #foreign keys per table
Semantic DatabaseEmbed Database embedding vector (768-dim)
ClusterID Topic ID from clustering (12 clusters)
CommunityID Community from graph structure (6133 communities)
Statistical DataVolume 12.3KB 215.9MB 1.5MB 73.7KB Total size of database files
AllJoinSize 10.00 $9\times 10^{21}$ $9\times 10^{16}$ 73.00 Row count when joining all tables
AverageCardinality 0.00 $7.3\times 10^{4}$ 67.56 15.03 Average distinct values per column
AverageSparsity 0.00 0.58 0.20 0.21 Average proportion of NULL values
AverageEntropy 0.60 12.20 3.50 3.28 Average entropy of columns
TABLE V: Summary of Edge (Database Relationship) Properties in WikiDBGraph ($\tau=0.94$)
Category Property Min Max Mean Median Description
Structural JaccardTable 0.00 1.00 0.12 0.02 Jaccard index of table name sets
JaccardColumn 0.00 1.00 0.36 0.37 Jaccard index of column name sets
JaccardType 0.33 1.00 0.93 1.00 Jaccard index of data type sets
HellingerDistanceType 0.00 0.57 0.07 0.02 Hellinger distance of data type distributions
GED 0 1230 111.89 4 Graph edit distance (GED) between schema graphs1
CommonTables 0 60 1.85 1 #common tables between schemas
CommonColumns 0 1092 74.00 13 #common columns between schemas
CommonDataTypes 2 6 3.24 2 #common data types between schemas
Semantic EmbedSim 0.9436 0.9998 0.9732 0.9771 Cosine similarity of database embeddings
SimilarityConfidence 0.9982 1.0000 0.9991 0.9991 Confidence score2 of similarity prediction
Statistical KLDivergence 0.00 34.01 17.05 17.60 KL divergence of shared column distributions
OverlapRatio 0.00 1.00 0.22 0.20 Ratio of overlapping values in shared columns
1The schema graph is a directed graph that connects tables (nodes) with foreign keys (edges), which is distinct from WikiDBGraph.
2The ratio of edges which has lower similarities.
VI Experiments
This section evaluates the utility of WikiDBGraph for collaborative learning. Section VI-A details the experimental settings. Because existing CL algorithms [55, 56] are not directly applicable end-to-end to WikiDBGraph, we design an automated data-mining pipeline (Section VI-C) that interfaces with existing CL training algorithms. The performance on the most similar database pairs are presented in Section VI-D.
VI-A Experimental Settings
This section describes the data, algorithms, licenses, and model/training hyperparameters used in our experiments.
Data. We select the top 2,000 database pairs with the highest similarity scores from WikiDBGraph and, among them, choose two representative pairs that have distinct TIDs yet exhibit high similarity in schema embeddings.
Algorithms. We evaluate representative collaborative learning algorithms spanning horizontal federated learning (HFL), vertical federated learning (VFL), and split learning. These algorithms include: Solo: Training on data from a single database. FedProx [28]: An HFL algorithm to mitigate data heterogeneity ($\mu=0.001$). FedOV [29]: An HFL method designed for data or feature overlap scenarios. SplitNN [10]: A split-learning algorithm for vertically partitioned data, where the model is split across clients. SFL [30]: A personalized federated learning method on a graph of clients. Combined: Centralized training on all clients’ data.
Model and Training Hyperparameters. The base model for all algorithms is a two-layer neural network with [64, 32] hidden dimensions. For SplitNN, this architecture is partitioned such that each client’s local model computes a 64-dimensional embedding, and a central server performs aggregation with the final 32-dimensional layer. All models are trained using the Adam optimizer [57] with a learning rate of $10^{-4}$ and a weight decay of $10^{-5}$.
Evaluation Metrics. We partition the data into an 80% training set and a 20% testing set. To provide a comprehensive performance comparison, we report four standard classification metrics: Accuracy, Precision, Recall, and F1-score. All reported results are the mean values over five independent runs on the combined test sets from both clients to ensure robustness.
Hardware. The experiments are conducted on a server equipped with AMD EPYC 9654 96-Core Processors (1.1 TB system RAM) and four NVIDIA H100 PCIe GPUs (80 GB HBM2e memory each).
Licenses. The WikiDBGraph dataset is released under the Creative Commons Attribution 4.0 International (CC BY 4.0) license. The code of WikiDBGraph is released under the Apache License 2.0.
VI-B Evaluation of Benchmark Design
This subsection evaluates the embedding-based relationship predictor against baselines and quantifies the contribution of each benchmark design choice via ablation studies.
Effect of prompt design. Table VI demonstrates the importance of prompting with both schema and data. The combined approach (Full) achieves the best overall performance, indicating that schema definitions and data samples offer complementary context.
Effect of encoder model. Table VII shows that the overall conclusions are consistent across encoder backbones and the string-based schema matching. Contrastive learning substantially improves AUC-ROC by 0.03–0.13 for different encoders. For instance, replacing BGE-M3 with all-mpnet-base-v2 [58] yields comparable performance after contrastive learning, suggesting that the gains primarily come from the proposed contrastive objective rather than a specific pretrained encoder.
Effect of negative samples. Fig. 5(a) shows that varying the number of negative samples has only a marginal impact on AUC. In particular, performance does not improve further beyond the default setting and slightly degrades when substantially increasing the number of negatives. We attribute this to the class imbalance in our data, where overly many negatives may bias the contrastive objective toward easy non-matches.
Effect of sample size in prompt. Fig. 5(b) shows that, once at least one tuple is included, the number of sampled tuples in the training prompt has only a limited impact on performance. Increasing the sample size primarily lengthens the context and raises computation cost, while yielding negligible improvements in AUC.
Effect of similarity threshold. Fig. 6 analyzes how the similarity threshold $\tau$ affects graph construction. As $\tau$ decreases, both the average node degree (Fig. 6(a)) and the total number of edges (Fig. 6(b)) grow rapidly, shifting the graph from sparse (dominated by highly similar pairs) to dense (containing many mid-similarity edges). Since most existing CL methods primarily work on highly similar pairs, we adopt a relatively high $\tau$ in our main setting for methodological compatibility. When more advanced CL techniques become available for mid-similarity regimes, our benchmark can readily support them by evaluating on pairs induced by lower thresholds.
TABLE VI: Performance of relationship prediction (optimal threshold, 5 seeds). Contrast.: contrastive learning. Full: Schema + Data
Configuration Method Accuracy AUC-ROC F1
BGE-M3 Original 0.9016±0.0026 0.9663±0.0012 0.9005±0.0029
Schema Only Contrast. 0.9622±0.0015 0.9908±0.0006 0.9618±0.0014
Data Only Contrast. 0.9766±0.0011 0.9964±0.0004 0.9765±0.0012
BGE-M3 + Full Contrast. 0.9772±0.0007 0.9967±0.0004 0.9770±0.0007
TABLE VII: Different encoder model’s test performance (optimal threshold, 5 seeds). Contrast.: contrastive learning.
Encoder Model Method AUC-ROC Precision Recall
String Match (Jaccard) - 0.7367±0.0051 0.7345±0.0117 0.5661±0.0025
BGE-M3 Original 0.9663±0.0012 0.9107±0.0037 0.8905±0.0066
Contrast. 0.9967±0.0004 0.9851±0.0029 0.9689±0.0027
all-mpnet-base-v2 Original 0.8667±0.0027 0.8369±0.0050 0.7401±0.0044
Contrast. 0.9968±0.0004 0.9846±0.0033 0.9827±0.0044
(a) #Negative samples
(b) #Sample size
Figure 5: Effect of hyperparameters on predictive AUC
(a) Averaged degree
(b) Total #edges
Figure 6: Effect of threshold $\tau$ on the graph properties
VI-C Automated Data Mining and Evaluation
We employ an automated data mining and evaluation pipeline to validate whether CL on database pairs identified by WikiDBGraph yields a performance improvement over Solo. This pipeline consists of five sequential stages. The detailed algorithm is presented in Algorithm 1.
The first stage, database pair sampling, selects the 2,000 pairs with the highest embedding similarity scores from all candidates that satisfy our validation criteria—both databases contain at least one table with a minimum of 100 rows. The second stage, label selection, identifies a target column for the learning task. We prioritize classification for simplicity; among all columns with a unique value count between 2 and 50, we select the one with the least number of categories to serve as the label. The third stage, table joining, constructs a comprehensive table for each database by left-joining its tables according to their foreign key relationships, starting from the table containing the selected label. To manage memory, this join process is halted if the table’s instance count exceeds one million. These heuristics enable a practical “zero-human-in-the-loop” pipeline, where the 2–50 value constraint filters out identifiers and continuous variables to ensure uniform classification tasks.
We then remove constant columns and any column with more than 50% missing values, followed by feature standardization. The fourth stage is column alignment, where we harmonize the schemas of the two resulting tables by retaining only their common columns. Column names are matched in a case-insensitive manner after ignoring all non-alphanumeric symbols. As an optional final quality-control step, we manually inspect each candidate task for semantic relatedness, column-alignment correctness, and label reasonableness. In total, 1,420 of 2,000 tasks pass and are used for training and evaluation. This step does not affect our conclusions, because the rejected tasks are dominated by invalid or non-executable cases already invalidated by the automatic pipeline (e.g., near-constant labels, too few non-constant features after joining/filtering, or runtime failures that prevent CL from training). Finally, the last stage is training and evaluation. For each client, the aligned data are partitioned into an 80%/20% train/test split. All CL algorithms are trained on the clients’ aligned training sets. To assess improvement, every algorithm—including Solo and Combined—is evaluated on the combined test set from all the clients.
In later stages, column alignment can be tackled via schema matching methods (e.g., [59, 60]) that explicitly recover semantic correspondences between columns. While [60] also employs a contrastive learning framework, it focuses on generating fine-grained contextualized column embeddings to resolve feature overlap, distinct from our work learning whole-database embeddings for global structural representation. Additionally, TURL [61] provides semantic table representations that facilitate table augmentation, further benefiting vertical alignment in instance-overlapped settings. Integrating these schema-level signals into future automated CL systems can significantly enhance overall integration performance.
Input: Database pairs with similarity scores; algorithm set $\mathcal{A}\subseteq\{\textit{FedAvg},\textit{FedProx},\textit{FedOV},\textit{SplitNN},\textit{Combined}\}$
Output: Trained models and performance of Solo and all $a\in\mathcal{A}$ on a combined test set
1
/* Stage 1: Database-pair sampling */
2 $\mathcal{P}\leftarrow$ select top 2,0002{,}000 database pairs with $\geq 100$ rows;
3 $\mathcal{R}\leftarrow\emptyset$;
4
5foreach pair $(D_{A},D_{B})\in\mathcal{P}$ do
  /* Stage 2--4: clean and align */
6  foreach $D\in\{D_{A},D_{B}\}$ do
7     $y_{D}\leftarrow\textbf{SelectLabel}(D)$;
     $T_{D}^{\mathrm{clean}}\leftarrow\textbf{JoinTables}(D,y_{D})$;
     // halt at 1M rows
8
9  end foreach
10 $(X_{A}^{\prime},X_{B}^{\prime})\leftarrow\textbf{AlignColumns}(T_{A}^{\mathrm{clean}},T_{B}^{\mathrm{clean}})$;
11  $X_{A}^{\mathrm{train}},X_{A}^{\mathrm{test}},X_{B}^{\mathrm{train}},X_{B}^{\mathrm{test}}\leftarrow$ partition $X_{A}^{\prime},X_{B}^{\prime}$ by 4:14{:}1;
12  $X^{\mathrm{test}}\leftarrow X_{A}^{\mathrm{test}}\cup X_{B}^{\mathrm{test}}$;
13
 /* Stage 5: train and evaluate */
14  $M^{\mathrm{Solo}}\leftarrow\textbf{Train}(\textit{Solo}\text{ on }X_{A}^{\mathrm{train}})$;
15  $\mathrm{Perf}^{\mathrm{Solo}}\leftarrow\textbf{Evaluate}(M^{\mathrm{Solo}}\text{ on }X^{\mathrm{test}})$;
16  foreach $a\in\mathcal{A}$ do
17     $M[a]\leftarrow\textbf{Train}\big(a\text{ on aligned }(X_{A}^{\mathrm{train}},X_{B}^{\mathrm{train}})\big)$;
18     $\mathrm{Perf}[a]\leftarrow\textbf{Evaluate}(M[a]\text{ on }X^{\mathrm{test}})$;
19
20  end foreach
21 $\mathcal{R}[(D_{A},D_{B})]\leftarrow\{\mathrm{Perf}^{\mathrm{Solo}},\mathrm{Perf}[\cdot]\}$;
22
23 end foreach
24
25return $\mathcal{R}$;
Algorithm 1 Automated Evaluation Pipeline
VI-D Performance of Existing CL Algorithms
The performance distributions of different CL algorithms are shown in Fig. 9. We report only the performance for the 1,207 that successfully completed training. Two primary observations emerge. First, the average scores of CL algorithms generally lie between the Solo and Combined baselines. Second, performance differences among CL algorithms are small; the notable exception is SCAFFOLD, which is known to be sensitive to hyperparameters and was not carefully tuned here. Both observations align with findings reported in NIIDBench [4], a well-known federated learning benchmark, indicating that evaluation using WikiDBGraph is effective and consistent with prior studies.
Despite the improvement in average performance, the generalization of existing CL algorithms to practical tasks—even when coupled with the automated data-mining pipeline (Alg. 1)—remains limited. Table VIII summarizes performance improvements across the 1,207 successful tasks: fewer than half of the tasks benefit from CL.
The high rate of non-improvement indicates that the primary bottleneck lies in data preprocessing. Our automated pipeline—which uses simple string-based column alignment—can miss semantic matches or, worse, join dissimilar columns. This flawed preparation yields a “garbage in, garbage out” effect: training on misaligned data degrades collaborative performance below the single-client baseline. Section VII supports this conclusion: manually curated alignments on similar database pairs produce substantial gains, implicating the pipeline rather than the CL algorithms.
While severe data heterogeneity (non-IID)—a known challenge for standard algorithms like FedAvg—is certainly a contributing factor, the stark performance gap between these automated results and the significant gains seen in our manually-curated case studies (Tables IX, XI) points to preprocessing as the more immediate and critical challenge.
Two major limitations of the current preprocessing pipeline are table joining and column alignment. First, the current implementation of the table-joining step is capped at one million rows to manage memory, which can discard valuable data; addressing this likely requires redesigning novel CL algorithms that take databases, rather than prejoined tables, as input. Second, the column-alignment step by comparing string similarity is straightforward but may be suboptimal because it ignores semantic similarity; addressing this calls for a new efficient, semantics-aware column-alignment method to determine which columns are equivalent. To validate this, we conduct an additional experiment using DeepJoin [62] for semantics-aware column alignment, while still using BGE-M3 as the base embedding model. As shown in Figure 7, on the same task set (with identical labels), semantic-aware alignment improves the mean F1 score by 0.06–0.09 for FedProx, FedOV, and FedAvg.
Figure 7: F1 improvement of DeepJoin for column alignment
Quantification of Non-IID Heterogeneity. To rigorously compare against synthetic non-IID benchmarks [29], we model the label skew as a multivariate Dirichlet process. We estimate the concentration parameter $\alpha$ using the robust Method of Moments [63], which aligns the theoretical Dirichlet variance with the empirical variance observed across WikiDBGraph pairs. The resulting distribution (Fig. 8) reveals a long-tailed spectrum of heterogeneity. While a substantial portion of pairs exhibit near-IID characteristics ($\alpha>5.0$), we identify a critical cluster of highly heterogeneous data: $24.7\%$ of pairs fall into the high-heterogeneity regime ($\alpha<1.0$), with $18.4\%$ exhibiting extreme non-IID skew ($\alpha<0.5$). This confirms that WikiDBGraph does not merely replicate standard distributions but captures the sharp, “long-tail” non-IID challenges inherent in real-world data silos.
Figure 8: Distribution of estimated Dirichlet-$\alpha$ in WikiDBGraph
(a) Distribution of F1-Score.
(b) Distribution of Accuracy.
Figure 9: Final metric distributions achieved by Solo, Combined, and different CL algorithms
TABLE VIII: Summary of automated data mining results with different CL algorithms on 2,000 database pairs
Metric Improved vs. Solo Total Ratio
FedAvg >> Solo 566 1207 46.9%
FedProx >> Solo 576 1207 47.7%
SCAFFOLD >> Solo 405 1207 33.6%
FedOV >> Solo 559 1207 46.3%
Combined >> Solo 565 1207 46.8%
VII Case Study
This section includes three case studies using manually curated data: feature overlap (Section VII-A), instance overlap (Section VII-B), and hybrid overlap (Section VII-D). All cases are novel findings enabled by WikiDBGraph, as their corresponding database pairs have distinct TIDs and no explicit links in Wikidata or WikiDBs.
VII-A Case 1: Two Databases with Feature Overlap
We exemplify the feature overlap scenario with a database pair: 02799 (Trypanosoma_Cruzi_Orthologs_Db_30, TID: Q62194121) and 79665 (TrypanosomaCruziOrthologs225, TID: Q62256692). This pair exhibits a maximum embedding similarity score (EmbedSim) of 0.9894. Both databases possess nearly identical schemas, each containing two tables with the same 24 columns and corresponding foreign key relationships, which constitutes the feature overlap. The primary distinction lies in their data instances and distributions; database 02799 contains 282 rows in both its tables, whereas database 79665 has 514 and 304 rows, respectively. This disparity makes the pair an ideal candidate for studying collaborative learning such as HFL. We therefore evaluate on a multi-class classification task to predict the label—details_encoded_protein with six distinct classes.
We first evaluate the performance of several neural network (NN) based HFL algorithms, with results summarized in Table IX. From these results, we draw two primary conclusions. First, all HFL methods consistently and significantly outperform at least one of the Solo baselines. This confirms the effectiveness of WikiDBGraph in finding new databases that enable effective collaborative training. Second, a notable performance gap persists between all HFL algorithms and the centralized training baseline (Combined). This behavior is characteristic of the challenges posed by non-identically and independently distributed (non-i.i.d.) data, a well-known issue in federated learning [4].
To demonstrate broader applicability, we extend our evaluation to tree-based models, known as FedTree [55] (Table X). The results consistently show that both the Combined and FedTree models outperform their respective Solo baselines. This indicates that the benefits of collaboration extend beyond NN, further validating the effectiveness of WikiDBGraph in identifying valuable opportunities for collaborative learning.
TABLE IX: Horizontal federated learning accuracy on feature-overlapped databases (six-class classification)
Method Accuracy Precision Recall F1-score
Solo (DB 02799) 0.2910 0.0980 0.2910 0.1443
Solo (DB 79665) 0.6187 0.4160 0.6187 0.4961
FedAvg 0.6154 0.6045 0.6154 0.5986
FedProx 0.6154 0.6045 0.6154 0.5986
FedOV 0.5753 0.5048 0.5753 0.5377
Combined 0.8997 0.8130 0.8997 0.8532
TABLE X: FedTree accuracy on feature overlap
Method Accuracy Precision Recall F1-Score
Solo (XGBoost DB1) 0.1720 0.1720 0.1720 0.1720
Solo (XGBoost DB2) 0.5071 0.5071 0.5071 0.5071
FedTree [55] 0.8428 0.8176 0.8428 0.8172
Combined (XGBoost) 0.9674 0.9674 0.9674 0.9674
VII-B Case 2: Two Databases with Instance Overlap
We demonstrate an instance overlap scenario with two databases of gene: 00381 (TrypanosomaCruziOrthologs1, TID: Q62194121) and 48804 (Ortholog_Lpg1l_Genomic_Data, TID: Q62256692), which exhibit a high embedding similarity (EmbedSim) of 0.95. Our analysis reveals that two tables, GeneOrthologsAnnotations from DB 00381 (the gene table) and Ortholog_Lpg1l_Protein_Annotations from DB 48804 (the protein table), share a significant number of instances based on their common GeneId column. We designate the protein table as the primary table for our task and the gene table as a secondary, feature-enriching table. We conduct a classification task that predicts the lpg_Uni_Prot_Protein_Id column from the protein table, which has 49 distinct categories.
TABLE XI: SplitNN accuracy on instance-overlapped databases (49-class classification)
Method Accuracy Precision Recall F1-score
Solo (DB 48804) 0.4082 0.3062 0.4082 0.3374
SplitNN 0.6531 0.5727 0.6531 0.5962
Combined 0.6782 0.6437 0.6782 0.6587
The results of NN-based algorithms, summarized in Table XI, yield two key findings. First, the split learning approach significantly outperforms the Solo baseline across all evaluation metrics. This result validates our hypothesis that enriching the feature set by incorporating data from the correlated gene table leads to substantial performance gains. Second, we observe a significant performance gap between SplitNN and Combined—the centralized baseline that represents the performance upper bound. While SplitNN captures the majority of the potential improvement, this gap highlights the inherent challenges of distributed model training compared to a centralized approach. In summary, these findings strongly underscore the practical value of leveraging instance-overlapped databases for collaborative learning evaluation.
Further experiments with a tree-based VFL algorithm, SecureBoost [56], are presented in Table XII, reinforcing our primary findings: across different algorithms, including specialized ones, extending the feature set consistently yields a significant performance boost. This demonstrates the general applicability of WikiDBGraph and further validates its effectiveness in identifying valuable pairs for feature enrichment.
TABLE XII: SecureBoost accuracy on instance overlap
Method Accuracy Precision Recall F1-Score
Solo 0.8418 0.8418 0.8418 0.8418
SecureBoost [56] (Combined) 0.9252 0.9252 0.9252 0.9252
VII-C Case 3: Multiple Databases with Instance Overlap
We consider a personalized CL classification task that reflects a policy analytics application. Each client corresponds to an independent country-profile database (DB10626, DB13841, DB17587, DB57240, DB73136) where features overlap only partially. The shared objective is to predict the minimum compulsory education age from locally stored attributes. To ensure the target is well-defined and comparable across databases, we discretize numeric ages into four ordinal classes (<5, 5, 6, 7). These five databases form a five-node, sixteen-edge subgraph, equipped with the WikiDBGraph node and edge properties. We then train a GNN-based CL method (SFL [30]) that consumes these graph properties; the results are reported in Table XIII.
TABLE XIII: SFL experiment results (per-DB averages over 5 seeds). Best in bold, second-best underlined. Gain (%) denotes the percentage of DBs where the method outperforms Solo, averaged over seeds.
Accuracy F1
[…]
Solo (No Federation) 0.8930±0.0636 N/A 0.8708±0.0671 N/A
FedAvg 0.9542±0.0444 20.0% 0.9373±0.0581 20.0%
SFL (No Properties) 0.9358±0.0509 16.0% 0.9138±0.0691 16.0%
SFL (Node Only) 0.9145±0.0756 16.0% 0.8979±0.0838 16.0%
SFL (Edge Only) 0.9559±0.0457 24.0% 0.9407±0.0606 24.0%
SFL (Both) 0.9559±0.0498 24.0% 0.9410±0.0655 24.0%
Table XIII supports two key findings. (1) Node and edge properties are critical for graph-based CL, enabling gains beyond FedAvg: enriching SFL [30] with node and/or edge attributes consistently improves both Accuracy and F1 over the no-property variant and yields the best overall performance, outperforming FedAvg. Notably, edge properties contribute the largest gains—the edge-only configuration nearly matches the full model—suggesting that pairwise cross-database signals are particularly informative for guiding message passing and personalization. (2) Graph-based CL improves over Solo, but structure alone is insufficient: SFL without any graph properties already improves over Solo, indicating that leveraging client connectivity is beneficial; however, it remains slightly below FedAvg, implying that the client graph structure by itself is not enough to beat federated learning baselines without informative node/edge properties.
VII-D Case 4: Multiple databases with Hybrid Overlap
To explore complex hybrid-overlap scenarios, we randomly select a small connected component (size 3 to 10) from WikiDBGraph for visualization (Fig. 1). The subgraph links a national monuments database (15418) to a cluster of historic-places registries (54379, 37176, 85770, 50469). Within the registry cluster, the databases are largely horizontally aligned, sharing many features such as AdministrativeEntity, AreaSize, and EstablishmentDate. The monuments database (15418) is further enriched by the national-treasure heritage registry (78243); these two databases share 542 columns (e.g., CategoryLabel, ArchitecturalStyle), indicating strong horizontal alignment. In contrast, 78243 and 85770 share only 32 columns—including identifiable attributes such as Latitude, Longitude, and AdministrativeEntity—and their instances do not fully overlap, yielding a partially vertical connection with limited instance overlap. The pair (78243, 85770) thus bridges two horizontally aligned clusters (national treasures and historic places), enabling collaborative tasks such as predicting the national-treasure category for historic places.
This component exemplifies a hybrid alignment pattern—simultaneous horizontal alignment within clusters and partial vertical alignment across clusters with incomplete instance overlap—that is not supported by existing CL algorithms. Consequently, graph-aware methods that operate over partially aligned, multi-database structures remain an open problem.
VIII Conclusion
This paper presented WikiDBGraph, a large-scale graph of 100,000 interconnected databases designed to address the critical lack of inter-database relationships in existing corpora. We demonstrated that WikiDBGraph successfully identifies correlated databases that benefit from collaborative learning, while also highlighting the performance gap compared to centralized models and surfacing novel challenges. WikiDBGraph thus serves as a crucial benchmark to measure this gap and guide future research into more effective and scalable collaborative learning paradigms.
AI-Generated Content Acknowledgement
GPT-5 is used for light polishing of human-written content. Claude and GPT-5 assist with code implementation. All contents are finally human-verified.
Acknowledgement
This research/project is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority.